### v2

# spacing=nonfrench

Abstract
This paper explores the use of Large Language Models (LLMs) in spear phishing message generation and evaluates their performance compared to human-authored counterparts. Our pilot study examines the effectiveness of smishing (SMS phishing) messages created by GPT-4 and human authors, which have been personalized for willing targets. The targets assessed these messages in a modified ranked-order experiment using a novel methodology we call TRAPD (Threshold Ranking Approach for Personalized Deception). Experiments involved ranking each spear phishing message from most to least convincing, providing qualitative feedback, and guessing which messages were human- or AI-generated. Results show that LLM-generated messages are often perceived as more convincing than those authored by humans, particularly job-related messages. Targets also struggled to distinguish between human- and AI-generated messages. We analyze different criteria the targets used to assess the persuasiveness and source of messages. This study aims to highlight the urgent need for further research and improved countermeasures against personalized AI-enabled social engineering attacks.
1 Introduction
Cybersecurity is a dynamic battle between attackers and defenders continuously adapting to new and emerging technologies. Although robust security measures can mitigate cyber risk, the strength of a system often depends on its users .[1] Among the myriad tactics exploited by threat actors, phishing remains the most common way to infiltrate systems. Phishing employs social engineering to persuade victims to perform an action, such as clicking on a malicious link or email attachment, that causes malicious code to run or disclose sensitive information. Phishing attacks usually take the form of email messages, telephone calls (vishing) or SMS messages (smishing) .[2] According to the FBI Internet Crime Complaint Center, phishing complaints in 2023 are five times more common than all other types of attacks, accounting for more than \$300 million in reported losses between 2021 and 2023 .[3]
Spear phishing is a targeted phishing method that uses personalized messages with specific information, such as a user’s name, job title, or home address, making them harder to distinguish from legitimate communications .[4] Targeted spear phishing poses a significant threat for the future as its prevalence and sophistication continue to grow. According to the Proofpoint "State of the Phish 2023" report, spear phishing accounted for 74% of phishing attacks in 2022, compared to only 8% for bulk phishing .[5] In instances where the pretext used against a victim matches their expectation, the attacker is likely to succeed .[6] For instance, users expecting a shipping update from a retailer are more likely to fall for phishing messages imitating such communications .[7] The effectiveness of spear phishing stems from its personalization, a characteristic that AI technologies increasingly excel at replicating, but with far less effort than required by humans.
Traditionally, crafting these targeted messages was less common. It required significant time and effort by cybercriminals to research and develop messages specifically tailored to their victims .[7] However, with the advancement of Artificial Intelligence (AI), emergent large language models (LLMs) can be used to carry out spear phishing campaigns with greater efficiency and human-like accuracy .[8] LLMs like GPT-4 generate text that mimics human language, making them potentially powerful in crafting deceptive messages with greater speed and less resources. The ability of AI to create these messages that are almost indistinguishable from human-written messages has heightened the need to assess how cybersecurity defenses can account for these new capabilities. While human-generated phishing may rely on individual creativity and context, AI-generated messages can leverage scalability and efficiency.
This study investigates the use of state-of-the-art language AI in crafting spear phishing SMS messages. We assess how current models fare in generating these personalized, malicious messages for distribution to potential targets and whether they currently perform on par or better than human creators. We also determine how well spear phishing targets can identify AI-generated messages, since such knowledge may be used to help protect oneself. Finally, we report qualitative feedback from targets on which factors of SMS spear phishing messages, such as personalization, tone, and word choice, contribute significantly to their effectiveness or failure. We do so with an eye toward the role of personalization, since it has the potential to both increase and decrease effectiveness.
The implications of this research are two-fold. First, awareness of the malicious capabilities of current technologies can equip cybersecurity professionals with knowledge and understanding on how to protect against the risks posed by AI-generated spear phishing attacks. Second, through a deeper understanding of themes and elements associated with more deceptive spear phishing attacks, cybersecurity education can evolve in tandem with emerging threats. Furthermore, distinguishing AI-generated messages from human-generated ones could shape future policy and training aimed at individuals to more reliably identify AI-generated content.
2 Review of Related Literature
The body of literature related to spear phishing is expanding, with numerous studies investigating the factors that lead individuals to fall for phishing attacks [6, 9, 1, 10], as well as the application of AI for detecting and generating such attacks .[8, 11] However, a notable gap exists in understanding how effectively AI-generated messages compare to human-authored ones, particularly in the context of spear phishing and SMS phishing (i.e., smishing), as discussed in this section.
2.1 Evolution of Phishing Techniques
As communication methods evolve, so do the techniques and methods used in phishing attacks. Early phishing attacks primarily relied on email due to its widespread use and ease of access to consumers and businesses .[1, 12] These emails generally used simple deception techniques, such as forged email headers, deceptive subject lines, and language that tries to copy legitimate and authoritative entities like banks and commercial platforms. Less successful phishing attacks were easily recognized by these identifying elements, as well as egregious spelling or grammatical errors. But over time, these attacks became more sophisticated, with attacks containing content that is almost indistinguishable from legitimate sources. Eventually, attackers began incorporating personalized elements to make these attacks even more convincing in what became known as spear phishing. These attacks use messages with some information about specific individuals extracted from social media, public records, or previous breaches, making them appear much more credible. The shift towards more targeted attacks enabled malicious actors to execute more successful phishing campaigns.[4, 5]
While email remains a dominant vector for phishing attacks, phishing campaigns have begun to adopt alternative platforms such as SMS text messages [13] and social media .[14] SMS phishing, also known as smishing, takes advantage of our growing reliance on text messaging via mobile devices. Smishing targets users with shorter, but equally deceptive text messages that often appear to come from trusted entities like delivery services, financial institutions or online platforms. Since many legitimate companies use SMS to automate user alerts, attackers may forge these messages to prompt users to urgently "resolve" a fabricated issue.
Social media platforms are also becoming fertile ground for phishing attacks. Attackers can create fake profiles of individuals or organizations to reach out to their targets or publish deceptive posts or ads .[14, 15] These can lead unsuspecting victims to open malicious links or give sensitive information. The open nature of SMS and social networks makes it difficult to combat phishing in these mediums. Despite the diversity of these attack vectors, existing studies rarely address the extent to which message origin—–AI or human—–contributes to their success or detection. This suggests a need to evaluate the current threat landscape.
2.2 AI-enabled Phishing
The integration of machine learning (ML) and neural networks into phishing attacks has opened new avenues for cybercriminals. With the development of large language models (LLMs), it has become easier to apply deep learning techniques to tackle tasks that demand high cognitive ability and creativity. In a study by Zhai et al., models like ChatGPT and GPT-4 were used to solve a variety of tasks, ranging from common-sense reasoning to more complex problem solving .[16] The AI consistently outperformed non-college students in these assessments, showcasing the potential for AI to execute creative tasks, such as social engineering.
Seymour and Tully trained a neural network on social media posts to create spear phishing messages to showcase the potential of AI to create targeted attacks.[11] Their approach reported a success rate of 30-66%, compared to the 5-14% reported from large-scale phishing campaigns .[17] Khan et al., utilized the GPT-2 model to generate phishing attacks, although models continue to improve dramatically over time .[18] Despite this, they used a game-theoretic approach to improve upon the model so as to gain a competitive edge by calculating the risks and rewards before implementing the attack. The study illustrated the potential to combine other disciplines to conduct more successful campaigns against potential targets.
As large language models (LLMs) have become more commonplace, continuous advancements in AI have significantly improved the capability of LLMs to generate human-like text. Models like GPT-4 (and GPT-4o), Gemini, and Llama have demonstrated an unprecedented ability to produce coherent, relevant, and persuasive content, which can pose significant cybersecurity challenges.[8, 19] The literature on AI-generated spear phishing is still emerging, with limited research specifically focusing on the cybersecurity risks associated with LLMs. There is a significant gap in understanding whether AI-generated messages are inherently more or less effective than human-generated ones and how defenders can distinguish between the two. This comparison has implications for detection technologies and end-user education.
AI-driven tools, like deepfakes [20] or automated reconnaissance [21], can significantly amplify the capabilities of AI-generated phishing. Further investigation is needed to evaluate how AI-enabled attacks can evolve in combination with other technologies and disciplines and what countermeasures may be effective in mitigating these risks.
2.3 Comparison of Human vs AI in Phishing
Comparing the effectiveness of human-generated versus AI-generated phishing attacks is a relatively unexplored area of research, yet one of growing importance. The success of human-generated spear phishing attacks traditionally relied on the skill of the person in crafting convincing messages tailored to the target. Each human attacker differs in their ability to gather data from available sources, identify vulnerabilities in reaching their target, and their capacity to use social engineering to their advantage. Koivisto and Grassini showed that the best human ideas still outperforms the capabilities of AI in high-level, divergent creative tasks.[22] No matter the skill level, crafting these highly personalized spear phishing campaigns are time consuming and resource intensive. Thus, the primary limitation of human-generated spear phishing is the lack of scalability, as each message must be individually crafted to its target to maximize the chance of success.
On the other hand, AI-generated phishing presents a solution to that shortcoming—where LLMs can generate phishing messages at scale, with significantly less human input .[8, 11] State-of-the-art models can quickly produce content that can mimic human natural language, which means that attackers can potentially execute highly targeted phishing campaigns with great efficiency. Currently, AI-generated phishing attacks can lack the contextual nuances that a skilled human author may provide, but when combined with principles from other fields [18] or new data analysis techniques[11], AI-enabled attacks could eventually make up for this limitation. Studies have yet to establish whether AI-generated phishing attacks can systematically outperform human-generated ones, particularly in real-world settings. Addressing this research gap is crucial to understanding how both AI and human-generated attacks exploit human vulnerabilities. The probability that AI evolves to outperform human counterparts to create more enticing spear phishing messages is high. Even now, the current iteration of LLMs poses a growing threat that requires new detection and prevention methods .[23]
3 Research Questions
We posit four research questions to investigate the capabilities of LLM generated spear phishing SMS messages compared to human capabilities. We hereafter refer to the research questions in this paper as RQ1, RQ2, RQ3 and RQ4:
  • RQ1.
    Are spear phishing SMS messages created by AI more convincing than those created by humans?
  • RQ2.
    What content characteristics contribute to a more convincing spear phishing message?
  • RQ3.
    Can people differentiate AI-generated spear phishing SMS messages from those generated by humans?
  • RQ4.
    What criteria do people use when identifying AI-generated spear phishing messages?
4 Methodology
This study comprises multiple steps using a novel methodology we call TRAPD, which stands for Threshold Ranking Approach for Personalized Deception. TRAPD is designed to analyze the perceived effectiveness of deceptive messages that have been personalized to a specific target (e.g., spear phishing messages) in an ethical manner. It uses quantitative methods to compare the perceived effectiveness of different types of messages (e.g., AI vs. human), as well as qualitative methods to understand what makes messages compelling (or not).
For this study, spear phishing SMS messages were created and evaluated by 24 human targets. Informed consent was received by the participants (i.e., targets), and the study was reviewed and approved by [anonymized IRB]. The following section describes the phases of the TRAPD Methodology, with details on how it was implemented in the current study to answer our research questions.
4.0 The TRAPD Methodology
One of the contributions of this paper is the introduction of the TRAPD methodology for evaluating personalized deceptive messages in an ethical manner. While the focus of this paper is on spear phishing SMS messages, the methodology may be used to evaluate other types of deceptive content tailored to individuals (e.g., personalized disinformation, other types of phishing messages). TRAPD is designed to facilitate comparison of different messages, such as those created by humans versus AI, or those created about different topics. Finally, it is designed to collect both quantitative and qualitative data, helping us to understand the "why" behind the quantitative findings.
At its core, the TRAPD methodology includes the following steps:
  1. 1.
    Recruit targets who willingly share personal information with potential attackers.
  2. 2.
    Generate personalized deceptive messages aimed at the targets (e.g., using humans or AI).
  3. 3.
    Have targets rank order (sort) the messages from most compelling to least compelling and choose a threshold above which they would be deceived.
  4. 4.
    Have targets provide qualitative assessments of their rationale for placing messages where they did.
  5. 5.
    (Optionally) Having targets “label" messages with a variable of interest (e.g., whether they believe a message was created by AI or not) and then provide qualitative explanations for their labeling choices.
Figure 1 illustrates the flow of the current project, as described by these TRAPD methodology steps. The following subsections explain how we implemented the TRAPD methodology in this particular study. We hope it will serve as a template for future studies that use the method to evaluate other types of personalized deceptive content.
Figure 1: Overview of the project within the context of the TRAPD Methodology steps. Recruitment, Generation and Assessment is numbered according to each step in the methodology.
4.1 Recruiting targets who share personal information
The first phase of research comprised recruiting participants as potential targets for spear phishing attacks and collecting personal information from them as part of a survey. Recruiting included convenience sampling from people whom the researchers know, placing printed fliers in the university library, posting online fliers on personal social media accounts, and sending online fliers to email distribution lists of university departments and neighborhoods. We sought to recruit targets with a variety of jobs and different ages, rather than simply having students participate. Fliers directed interested individuals to complete a Qualtrics survey to sign up, consent to participate, and share some personal information for the purposes of this study. Participants were also told they would be invited to participate in an in-person interview where they would evaluate personalized messages created for them based on the information they provided. Participants were not told at recruitment time how the personalized messages would be generated (i.e., human vs AI).
A total of 41 people (i.e., targets) completely filled out the survey, which asked for their name, email address (to schedule the follow-up interview), sex, age group, and self-reported technical ability. Targets also provided details about themselves that would be used to generate the spear phishing messages including: participant hobbies, city and state in which they live, their occupation and workplace, and something from their home that they recently posted on social media.
4.2 Creating personalized deceptive messages
Using the personal data gathered from each participant, spear phishing messages were generated from both the GPT-4 model and from human authors. The goal of each message was to convince the human target to click the link included in each message. The spear phishing messages that were generated were in the form of SMS messages. Each prompt focused on one of three topics: the target’s hobbies, workplace, or something they own that they recently posted on social media. Below is an example of the prompt for one of the targets (name altered, as it was in the original prompts to maintain target’s anonymity):
Make a spearphishing message in the form of an SMS for Ethan Clark, a 18-25 year old Male who recently shared about their pepper plants; they would post pictures of how they are flowering and growing peppers. Have them click a link that is not from a legitimate source, but is believable.
The same exact prompt was provided to human authors and GPT-4 as described in the following sections.
4.2.1 Human Generation
Human authors were recruited from undergraduate students enrolled in a university cybersecurity program or honors students taking a course on deception during Winter semester 2023. All students had learned about techniques used in phishing attacks and had practiced creating them as part of a class. They were given 15 minutes of time during classes to write up to 4 messages, each based on a different randomly selected user and prompt. Students received extra credit for their participation. Based on observations, students were highly invested in the task as evidenced by a lack of discussion or participation in other activities. A few stayed longer than the 15 minutes to finish their messages.
Authors typed their spear phishing messages into an online survey tool that allowed for emojis and had a text box that approximated the size/length of a standard SMS message. They were provided with 4 randomly selected prompts containing information about their target, then asked to generate an SMS spear phishing message that would use that information in a way that would get the user to click on a link. After each prompt, the author was asked to assess their confidence in the ability of the message to deceive the target. They also answered questions about their age group, sex, technical ability, and understanding and experience relative to spear phishing.
Ninety-nine student authors participated, with the intent of creating a balanced pool of message contributions across demographics. In total, 363 messages were gathered and screened for invalid entries that were not spear phishing messages or were obviously problematic (e.g., not completed). In cases where a user included only a placeholder for a link (e.g., "[link]") we replaced with a generic tinyurl link.
We expected that some human authors may not have created valid messages. For this reason, we collected at least 3 messages for each target/topic combination (e.g., target15/hobbie), even though we ultimately only showed targets 2 messages per topic. To give the human authors the best chance, all messages underwent an additional review by our team, which include two cybersecurity professors that teach courses that cover social engineering. We excluded any messages that were incomplete, did not include a prompt to click a link, or were of obviously low quality. In cases where there were more than two valid messages for each topic (work, hobbies, social) for each target, two were selected at random to proceed to the next phase. All in all, 246 human-authored messages were retained and prepared for the next phase. This process helped counter potential problems of having students be the authors of spear phishing messages.
4.2.2 AI Generation
A script was written to call the GPT-4 API to generate spear phishing message outputs automatically. The same three prompts given to the student authors for each target were fed to the script to generate AI-generated spear phishing messages. An example of a GPT-4 output is shown below:
Hey Stephanie! I came across this incredible home organization app that I think will help you streamline your daily tasks and save time. It’s been a game-changer for me! They are currently giving away a free 1-year subscription to the first 100 users who sign up. Don’t miss out on this opportunity! Here’s the link: bit.ly/organizehome4u
Stay organized!
Your friend :)
Since GPT-4 outputs are not completely reproducible, three different responses were gathered from each prompt. This resulted in 9 messages generated by the model for each target. Similar to what was done with the human-authored messages, we selected at random 2 messages for each topic, totaling 246 AI-generated messages to be used for the next phase.
After both sub-phases were completed, each target had 12 spear phishing messages: 6 created by GPT-4 (2 for each of the 3 topics) and 6 created by humans (2 for each of the 3 topics). By the end of this phase, 492 simulated spear phishing “attacks” were prepared for assessment.
4.3 Target Interview and Sorting Activity
All 41 people who filled out the survey were invited to come back for the interview/sorting activity approximately a month later. Only 25 (61%) responded and arrived on their scheduled interview day. The distribution of participant demographics are shown on Figure 2. We were successful in recruiting a range of ages, as well as professions. Although 15 (60%) of the targets were affiliated with the university, only 8 (32%) were students. A range of jobs were reported (e.g., librarian, software engineer, instructional designer, sales agent, teaching assistant).
Figure 2: Target Demographics (n=25).
4.3.1 Threshold rank order
The 12 spear phishing SMS messages were shown to each target during the interview, with all messages displayed on a separate printed piece of paper showing a phone and text message image around the content. Participants were asked to arrange messages in order of most to least likely to lead to the target clicking on them. Figure 3 shows an example of how the targets laid out the messages in the experiment, illustrating the ranked order from most to least convincing, as well as the threshold where they would have started to click the link based on the participants’ assessments. Participants were not explicitly told the messages were spear phishing messages when they were asked to sort them.
Figure 3: A sample of the threshold rank ordering used in the target interview. After ranking the messages from most (red) to least (blue) likely, a marker (yellow) is placed to indicate the threshold for clicking.
4.3.2 Qualitative Assessment
Participants were then asked what elements of the messages led them to be tricked, as well as other feedback that they were willing to share regarding the criteria they used to sort messages. They were also asked what made them place the less convincing messages at the bottom. They were encouraged to look at individual messages when providing feedback in order to ground their comments with specific examples. This was often followed by more general assessments by targets about overall criteria they used that applied to more than a single message.
4.3.3 Label as Human vs AI
Participants were then told that "one or more of the messages were created by an AI," and were asked to place an AI "marker" on the message or messages they believe were created by AI as shown in Figure 4. They were then asked what made them think the messages they chose were made by humans or AI. Again, they were asked to use specific messages to ground their observations, followed by their overall thoughts. At the target’s request, researchers revealed which messages were made by AI vs human authors.
Audio recordings were made for all 25 interviews which were then assimilated for transcription and further analysis. Images of all sorting and labeling were also captured for data analysis.
Figure 4: An example on the AI labeling phase of the target interview.
4.4 Statistical Methodology
Several statistical analyses were conducted to evaluate the difference in performance of AI-generated and human-authored spear phishing messages (RQ1), the difference in performance of particular elements in the spear phishing messages (RQ2), the overall accuracy of the targets in identifying AI-generated messages (RQ3), and whether certain factors in the messages helped them identify AI-generated messages better (RQ4). All statistical analyses were carried out in R.
To address RQ1, we compared the average ranking of both types of messages across all the targets. To verify statistical significance, we utilized a permutation test [24] to validate our findings. Following this, we applied a logistic regression model [25] to predict the likelihood of a subject saying they would click on a link (i.e., being above the click threshold) based on whether the message was AI or human. To account for potential correlations within subjects and authors, we fitted a generalized linear mixed-effects model (GLMM).[26] To assess how similar or better the performance of AI messages are compared to human messages, we also employed a Bayesian logistic model[27], using a horseshoe prior on the log odds ratio, pulling estimates towards zero when evidence was weak but not excessively when evidence was strong.
To address RQ2, we compared the average rankings of the messages according to each of the three topics: Job, Hobby or Social. We also used a logistic regression model based on the threshold self-assessments in relation to the three topics to assess whether a topic was more successful in generating clicks than the others.
To address RQ3, we tallied the correct and incorrect predictions for all targets across AI and human messages. We also developed a predictive model to assess whether subjects could correctly identify the origin of the messages based on factors (RQ4) such as the presence of emojis (Emoji), modifications to links (LinkMod), and the total number of characters in the message (CharacterCount). This is to determine if these features could significantly aid subjects in distinguishing between AI and human messages.
5 Results
This section describes the statistical and qualitative results that address our research questions.
5.1 Statistical Analysis
The main metrics for performance of each topic (job, hobby and social) and source (Human and AI) is their average ranking as determined by the order which their corresponding messages were placed by the targets, and the click probability, as determined by the threshold marker. Table 1 shows a summary of the average ranking and click probability of each topic and source. This subsection expounds on the statistical significance of each result.
Topic Ave. Rank Click Rate
Job 5.71 38%
Hobby 6.66 19%
Social 7.13 17%
Source
Human 6.58 21.33%
AI 6.41 28%
Table 1: Average Rank and Click Rate by Topic and Source. Job-related and AI-generated messages generated more clicks than their counterparts.
5.1.1 AI vs. Human Ranking and Click Probability
To assess performance between AI and human messages (RQ1), we compared their performance based on how high they were ranked in the target interviews and how often each type of message would be clicked by our targets. On average, AI-generated messages ranked slightly higher (6.41) than human-authored messages (6.58). Figure 5 describes the distribution of the AI and human messages across each rank.
Figure 5: Rank Distribution from each source. On average, AI-generated messages ranked slightly higher than human-generated messages.
Results from our statistical analysis did not indicate significant differences between the two groups, with a p-value of 0.665. To account for the non-normal distribution of ranks and the variation across targets, a permutation test was conducted, but it again showed no significant difference in ranks between AI and human-generated messages.
The logistic regression model came with similar results. The model did not find a significant difference, with a p-value of 0.182. The predicted probabilities of clicking were 21.3% for human-generated messages and 28.0% for AI-generated messages. Although the AI-generated messages had a higher predicted click rate, this difference was not statistically significant. Fitting the GLMM also yielded similar results, indicating greater variation between subjects than between authors.
The odds ratio for clicking on a link in an AI-generated message versus a human-authored one was 1.43, with a 95% confidence interval of 0.847 to 2.446. This wide interval includes 1, indicating that the increased likelihood (43%) of clicking on AI-generated messages compared to human-authored ones is not statistically significant and could range from 15% lower to 145% higher.
We employed a Bayesian logistic model to further the analysis. Using a horseshoe prior on the log odds ratio, we obtained a posterior mean odds ratio of 1.22 (95% CI: 0.86 to 2.05) with a median posterior of 1.16. From this posterior distribution, we estimated that the probability of the odds ratio being less than 1 (indicating humans are better than AI) is 19.7%, while the probability of it being greater than 1 (indicating AI is better than humans) is 80.3%. The probability of the odds ratio exceeding 1.5 (AI being more than 50% better than humans at generating clicks) is 19.9%. A plot of the posterior distribution of the odds ratio shown in Figure 6 indicates that, while there is a long right tail suggesting a possibility of AI being significantly better, most of the probability mass is near 1, suggesting neither being definitively better.
Figure 6: Posterior distribution of the odds ratio of Human vs AI messages. Results suggest that AI-generated messages performed as well if not better than human-authored messages.
5.1.2 Topic-based Ranking & Click Probability
One of the key content characteristics of the messages (RQ2) relates to their topic: Job, Hobby, or Social media post. On average, Job ranked the highest (5.71), followed by Hobby (6.66) with Social ranking the lowest (7.13). The distribution of these rankings is shown in Figure 7. Using a Tukey post-hoc test to analyze statistical significance of the ranking of each topic, we derived the confidence intervals of each pair of topics. Results show that the mean rank for Job is significantly higher than Social, while Job-Hobby is not statistically significant, and much less for Hobby-Social.
Figure 7: Rank Distribution of Each Topic. Job-related messages ranked higher than hobby and social-related messages.
We used a logistic regression model to determine if the topic of the message influenced the likelihood of clicking. The model showed an overall p-value of 0.0009, suggesting that the topic played a significant role in generating clicks. Job-related messages showed the highest probability of clicking (38%) while Hobby-related (19%) and Social-related (17%) messages scored much lower. Pairwise comparisons showed that Job had a significantly higher click probability compared to Hobby and Social, with no significant difference between Hobby and Social topics. Specifically, the comparison between Job and Hobby had an estimate of 0.9605 (95% CI: 0.1933 to 1.7277), while Job vs Social had an estimate of 1.0961 (95% CI: 0.3081 to 1.8841).
5.1.3 Identifying Message Origin
To assess if subjects could correctly identify whether a message was AI-generated or human-authored (RQ3), we analyzed their guesses for both types of messages. Subjects correctly identified the message origin 52% of the time. Note that 50% would be expected from randomly guessing. Table 2 describes the confusion matrix that details their accuracy.
AI Human SUM
AI (truth) 78 72 150
Human (truth) 72 78 150
SUM 150 (52%) 150 (52%) 156 (52%)
Table 2: AI Identification Confusion Matrix. Results suggest that target accuracy is not much better than flipping a coin.
We also used a logistic regression model incorporating predictors such as the presence of emojis (Emoji), whether there were modified links (LinkMod), and the number of characters in the message (CharacterCount). The overall p-value of 0.3253 indicated no significant ability to distinguish between AI and human-authored messages based on these features.
5.1.4 Demographics and Phishing Experience
Participants in this study were roughly half male and half female and had various occupational roles including entry-level employees, mid-level managers, and college students. Despite this diversity, no statistically significant differences were observed across demographic groups for click rates or the ability to identify message origin (AI vs. human). Prior experience with phishing did not appear to significantly improve performance in identifying AI-generated messages or reducing click rates, aligning with prior research suggesting that familiarity alone may not translate to better detection.[28]
5.2 Content Characteristics of Persuasive Messages
This section addresses RQ2 from a qualitative perspective. Specifically, the following subsections identify the core themes that resulted from our analysis of the verbal explanations participants gave when describing why they sorted messages as most likely to deceive them or least likely to deceive them. Table 3 provides a summary of the percentage of targets who mentioned these themes that influence their perception of a convincing message.
Themes % Mentioned
Relevance 76%
URL 64%
Medium 40%
Style 40%
Scarcity 32%
Inaccuracies 28%
Rewards 28%
Sender 28%
Table 3: Key Content Characteristics Influencing Message Persuasiveness. Most messages that were convincing to targets were identified to be "relevant" to them to some degree.
5.2.1 Personal Relevance
The messages being personally relevant (or irrelevant) to the participants was mentioned in 76% (19 of 25) of the interviews. Seventeen targets explained that personal relevance of the message made it more convincing. Several participants pointed to messages directly tied to their occupational responsibilities, with one saying that “because that is my job is to help people with records…this is one that I feel like I’m most likely to engage with" (T13). Another target, who is a banker at a credit union, noted the similarities between the targeted message and messages s/he receives at work. S/he noted, the “alert literally looks like the alert we get [at work] when there’s a fraud” (T33). Some participants pointed out how the spear phishing messages aligned with their personal interests. One expressed enthusiasm about a message describing a “Vineyard gardening club”, believing that “somebody from our community” may have sent it. They expressed that they “would love to get involved in that" (T31). While most participants were more convinced when messages were relevant to them, one more skeptical participant mentioned that although “this interests me and it could exist, but I’m not gonna click on it” (T08).
On the other hand, personal relevance was perceived as unconvincing to 17 of the targets. Some participants highlighted messages that didn’t align with their current activities. For example, one interviewee noted, “I’m also not actively dancing anymore, so that’s just weird that they’re offering me something like this” (T25). Another target was suspicious of a social media-related message because they “don’t have an Instagram [account]” (T34). Some targets emphasized messages that failed to pique their interest, with one saying that they are “not into Brandon Sanderson [the author]” (T37). Another mentioned that they have already rescued an animal and “don’t need more right now” (T02). Some targets mentioned messages that contained information that they had not disclosed. One target noted that the message wasn’t “legit” because they “never put [their] studio equipment online” (T31), while another mentioned that “it’s a little weird to be selected for something that you don’t apply for” (T16). Participants also raised suspicion about messages related to unfamiliar organizations or activities, mentioning that they will not apply because “I don’t know that organization” (T41), or they “…have no idea what that would be about. So I’m just going to ignore it” (T08). Finally, targets showed skepticism when messages were somewhat relevant but lacked specificity. One target noted that “there just wasn’t enough content in the message body that was specifically directed at me” (T20). Another target highlighted a message about “recent health challenges” but noted that “there’s no indication of like, who is supposedly sending the message” (T18). Participants described how personal relevance within the message provides a very strong indicator for the message being perceived as convincing or not.
5.2.2 URL
The URL was mentioned by 64% of all participants as having convincing attributes (16/25). Six participants noted how the URL can provide a sense of trust, with one noting, “[my university], I kind of trust [my university] and the URL here" (T41). Three participants mentioned convincing characteristics with regards to the domain of the URL. One stated, “The biggest part of this one [URL], and I even debated whether I would put it first or second; but it has an .edu link . . . and maybe I’m very ignorant about this, but I feel like that’s harder to, like, create a fake .edu website” (T16). In contrast, another participant used similar reasoning for not clicking on a link, noting, “At first, I thought, Oh, I would click on that. But then it says google.com.org; which is weird, because I’ve never seen that before” (T07). This same participant mentioned some inaccuracies related to the domain that would cause him/her to rethink clicking on it. Another participant mentioned the use of a .net website which contributed to them clicking on the link because the domain seemed trustworthy.
Two participants were more likely to click on the link if it included the HTTPS protocol. One said that “If it looks legit, yeah, like HTTPS and the /ProvoLibrary, something like that” (T41), then they would click. Another participant said, “they all say HTTPS, which makes me feel like it’s secure” (T07). Personal association with a URL may also play a role, with one participant stating s/he was more likely to click on the link if it included his name within the URL saying that “the link actually has my name in it. That would have really like, definitely thrown me for a loop, like really made me like, actually think about clicking the link just because it’s like, oh, this is actually from maybe from [my university]” (T20).
Ten participants were concerned about the use of a link shortener within the message. One mentioned that “I feel like anytime I’ve seen a tiny URL address it, like, either hasn’t been real, or it’s been like, weird or just different things” (T16). Six of the participants were less likely to click on a link if it was misspelled e.g., “facebock” instead of “facebook” (T13). Overall, the perception is that URL name, domain, and indicators within the domain (e.g., HTTPS, spelling) help create perceptions about the deceitful nature of the message.
5.2.3 Technology Communication Medium
The technology medium used for communication was mentioned by 40% (10/25) of the interviewees. One of the 10 mentioned the medium increasing their likelihood of clicking on the link. S/he mentioned that it is “not that abnormal for [them] to receive” text messages from their home city" (T13). However, most targets saw the text medium as a red flag, even when they were inclined to believe the content. This was sometimes because of warnings from trainings they had received. One explained that “at work we were warned that all, like, valid messages will be through emails” (T25) and “This one, [employer] is not going to reach out to me via text for security stuff, period” (T08). T37 mentioned that they don’t have a “company texting deal,” making them more suspicious of SMS messages. Another men-tioned that “it’s very rare that people message me, you know, we talk over Slack or we talk over WhatsApp, maybe through email” (T31). Others noted that other communication tools, such as an online learning platform (T09) or a genealogy website (T02), would not have sent them a text message. Interestingly, in all these cases, the technology communications medium was associated with the source of the message. In other words, receiving a text message was not abnormal in and of itself, but receiving a text message from a particular source was deemed either appropriate or not.
5.2.4 Messaging Style
Ten respondents (40%) mentioned issues related to message styling, such as text formatting, tone, and structure. Four participants acknowledged that style played a role in enhancing the credibility of a message. For example, T09 affirmed that a message was “more convincing” because the formatting and tone felt “kind of personal.” T31 also emphasized formatting in relation to message personalization explaining, “First, they talk about how they found me. They loved my work. And they saw my music profile. They talk about a music festival…. The format seems like it’s from an organized festival.” A message’s “casual style” was also suggested by respondents as a way for the message to feel authentic. A message’s “less formal” style was more “enticing” and “attractive” to T41. On the contrary, two participants mentioned that the message tone was off-putting. For example, T09 identified a message as potentially a phishing attack due to its “sales-y” tone. Additionally, T18 mentioned how sales-oriented “buzzwords” indicated a message “feels more phishy.”
A quarter of respondents (6 of 25) discussed the role of emojis in the context of message style. All six asserted that emojis diminished the credibility of the message. T34 expressed that “I think a lot of emojis is kind of something that usually is a red flag for me.” Multiple respondents conveyed that they did not anticipate receiving emoji-filled messages from senders claiming to represent professional organizations. For instance, T27 said that in the case of a legitimate organization such as a city biking club, emojis would be “out of place." Similarly, T20 explained that if a message claimed to be from their university, the presence of emojis would cause them to feel “thrown off." From the perspective of participants, the use of emojis in messages, particularly those seemingly sent by professional groups, caused such messages to lose legitimacy.
5.2.5 Scarcity Principle
Urgent wording was mentioned by 32% of participants (8 of 25). Two of the eight mentioned they were more likely to click on a link if it included urgent or fear inciting language. One explained, “And because of ’suspicious activity’ under my university account, then my first thing is, well, I better click on this. Because if there’s an issue with my account, I want to fix it right away” (T34). The other target mentioned the same thing, that they would click on the link because they wanted to fix the problem as soon as possible. They explained, “And just like, I’d read through it real quick, and just click on it because I’m like, oh, shoot, that’s maybe something important that’s going on. Okay”(T06).
The remaining participants (6 of 25) said they were less likely to click on the link if it included tactics such as fear and urgency. These participants explained that time sensitive or urgent messages feel like a warning sign. “That’s always a way of like, okay, they’re trying to get you to shut off your logical brain. Put yourself into ’Oh, my gosh, we have to do this right now.’ Yeah. Right.” Another participant confirmed, “If it seems urgent to you, you’re not going to click it because it’s very dangerous” (T13). Similarly, T41 explained, “I don’t like anything where it’s like hurry, offer and send because that always makes me wary” (T07). Similar to the their concerns with time-sensitive messages, participants also discussed how fear inciting messages can cause them to logically analyze the message at a deeper level. One explained, “To the message itself, I was a little bit unsure about just like, the fact that it was so time sensitive, and that it was like, kind of threatening to lock me out of all of my educational accounts if I didn’t do what it asked for. Because that’s not typically what I’ve seen be the case” (T16). And another stated, “Yeah. And it’s the idea of a new virus for young dogs. I’m going that’s fear mongering” (T01). The scarcity principle has long been used in literature to deceive readers. For the spear phishing messages in this study, participants had mixed reactions. Some felt the intended pressure to click, while others saw the language as manipulative and as a warning sign marking potential danger.
5.2.6 Context Inaccuracies
The presence of inaccuracies in the messages was mentioned by 28% (7 of 25) of the targets. These inaccuracies contributed to perceptions about message credibility. Participants identified discrepancies related to the names or characteristics of individuals or enti-ties within their professional spheres. For example, two targets remarked about the message mentioning a colleague that does not exist, stating “there’s no Mike at work” (T37) or “there’s no one named Sarah on the [my university] instructional design team” (T18). Another participant expressed confusion about being questioned on matters not relevant to their professional duties, saying that they “don’t deal with payables. So why is that asking me regarding payment?” (T25). Another target noted discrepencies related to job responsibilities, asserting that such library-related decisions would “go through me at the library”, making any logistical deviation “immediately sus-picious” (T01). Regarding messages related to their hobbies, one participant raised skepticism about the message claiming a “great deal for a hiking trail” which the participant knows through experience doesn’t actually require payment for use (T30). Another target was also suspicious of the message being from a library that they “just don’t use… so they would not know me to contact me” (T08). In summary, the instances where inaccuracies were mentioned ranged from misrepresented personnel and responsibilities to factual discrepancies about the targets’ affiliations, activities, and expectations.
5.2.7 Plausible Rewards
Realistic rewards significantly influenced the perceived credibility of phishing messages, as noted by 28% of the participants (7 out of 25). One participant said they would be more likely to click on a link, “because the reward is connected to something I put so much time and effort into" (T31), while another participant stressed the persuasiveness of plausible or realistic offers, mentioning, “It’s like taking you to a link to look at an offer, but it’s not a crazy insane offer" (T16). Some targets expressed enthusiasm for realistic opportunities, stating, “It’s something I’ll be really excited about, but also something that I think is realistic" (T34). Conversely, 24% of participants (6 out of 25) identified messages with rewards that seemed “too good to be true" as less convincing. One participant mentioned it’s “probably not likely” to be the “developer of the month four times a year” (T37). One of the targets questioned the credibility of messages offering free access to goods and services, stating, “It sounds too good to be true" (T34). Another target warned against messages promoting anything for free, noting, “Anything that’s free is a little too good to be true. Okay, so that’s when I would be very careful" (T27). This collective sentiment emphasized the convincing nature of realistic rewards in phishing messages while avoiding extravagant or implausible claims that trigger suspicion.
5.2.8 Sender Familiarity
Participants described how the perceived credibility of phishing messages is tied to the sender’s identity, as noted by 7 out of 25 targets (28%). One participant said the message seemed more convincing because the sender “introduced herself as [their] neighbor" (T30). Another explained the importance of a personal introduction, stating, “The sender started with ’My name is John’" (T41), which helped the target to trust the message. Several participants expressed the importance of connecting with the senders of the messages. One target was convinced by the “actual company branding” that mimicked promotional texts they had previously received (T18). Another explained, “This could be someone from my ward [church congregation], wanting me to like check out some plants or something" (T31). Participants explained that when the message sender appears to be a legitimate, familiar source, messages seemed more credible.
Shifting towards participants’ suspicions, a substantial 68% (17 out of 25) highlighted the pivotal role of the sender in raising skepticism about received messages. One participant explained that messages appear suspicious when the introduction is not consistent with other messages from that same sender, saying, “James is my boss… it’d be weird that he was introducing himself that way… that doesn’t sound like my boss" (T31). Several participants described the need to understand how the sender unexpectedly obtained their phone number and the purpose for which the sender sent a message. One target noted, “So like, how did you get my number in the first place?" (T27), while another questioned, “Like why would Lowe’s have my phone number?" (T20). Participants described how recognition of the sender’s phone number is also an important component of message believability. One participant explained, “If it’s from a number I don’t recognize at all. Then yeah, probably the first thing I would ask is who it is. And then however, if the person is like, if their name is like the same as one of my friends, then I probably would open it" (T34). Participants described that messages lacking personal connections from the sender raised suspicions, with one stating, “I don’t feel like it’s anyone who’s personally connected with me; it just seems like spam, you know?" (T31). Being familiar with the message sender was described as a strong indicator in whether subjects believed the messages to be legitimate. It is important to note that none of the study text messages sent to participants indicated a source phone number. Rather, our focus was on the content of the messages.
5.3 Human/AI Source Identification
This subsection addresses RQ3, which asks how effective targets are at identifying AI-generated spear phishing SMS messages. The statistical analysis presented earlier showed that targets could not effectively identify AI-generated messages. The qualitative results presented here reinforce the difficulty of this task for targets. Although the targets described some of their reasoning behind their guesses on which messages are AI-generated (see following section), 12 out of 25 of the participants stated in some way that they were uncertain about their decisions, often relying on intuition rather than any specific criteria. One target remarked, “sometimes it’s a gut feeling maybe more than like a specific thing you’re looking for?" (T07), while another said it’s “just a feeling" (T02). Others attributed their lack of criteria to advancements in AI, saying that they,“don’t really, really know my criteria, because I know that AI is getting so good" (T04). Several targets openly admitted their lack of expertise, saying they “have no idea" (T15), and “actually I don’t know, I don’t know" (T23). Some were unsure if AI could personalize content or include icons in text (T02), while others conveyed concerns of AI’s competency, hoping that “the ones that are worse would hopefully be the AI ones" (T34). Overall, these responses show the difficulty the targets faced in distinguishing between AI and human-generated content. In general, they did not seem to have mental models that provided meaningful guidance in identifying AI-generated messages.
5.4 AI Identification Criteria
This subsection addresses RQ4 by identifying the criteria people use to try and identify AI-generated messages. Table 4 provides a summary of the percentage of targets who mentioned these criteria (or lack thereof) that influenced their determination of an AI-generated message.
Themes % Mentioned
No idea 48%
Style 40%
Emojis 20%
Personalization 32%
Word Choice 24%
Structure 24%
Grammar 24%
Message Length 16%
URL 8%
Table 4: Themes Influencing Perception of Messages as Human or AI-Generated. Although targets shared differing criteria for identifying the source of the message, a significant number do not have an idea.
5.4.1 Style
The message style was mentioned by 40% (10/25) of the targets. Many targets noted that AI-generated messages tend to be, “pretty formal" (T21) or “overly informal or overly formal" (T09) while others said messages often appear “extremely generic sounding," as if they were created from a prompt with specific parameters (T18). Additionally, the presence of “glaring flaws" such as typos and awkward phrasings were attributed to being human-made mistakes. Such common mistakes contrasted with the “more robotic" nature of AI messages (T08). Some targets observed that AI-generated messages might seem “too perfect," such as refraining from casual texting language that humans often use, such as “RN" for “right now" (T34). One participant noted that the excessive use of exclamation points is what made them think the message is AI generated: “I mean, honestly, what’s driving me insane about all of these is exclamation points. Yeah. I’m like, why are there so many exclamation points all over?" (T07). In contrast, messages that gave a sense of urgency were more likely to be perceived as human-written, as one target explained: “Then I didn’t put this one because they were actually pressuring me to do it immediately. So I don’t think AI can do it" (T02). Overall, targets indicated that a message’s style played a role in their AI vs human generated message decisions, with AI messages being attributed to using a style that is overly formal, generic, or laden with exclamation points; and with human messages appearing more casual, urgent, and imperfect.
5.4.2 Emojis
The use of emojis in messages was mentioned by 20% (5/25) of the targets as being indicative of an AI-generated message. Two concluded that the use of emojis makes a message appear to be written by AI (T25), with one target explaining, “there’s always an emoji” and that the emoji messages don’t look very personal (T30), indicating a machine wrote the message. Alternatively, three of the targets connected emoji use to human writing because they were not sure if AI could use emojis. One target claimed they didn’t think a message was AI written because of “the icon” (T02), while another explained “I’ve never asked an AI to do something with emojis. I wasn’t sure if it could. I’ve never seen that, but I wouldn’t expect emojis” (T18). The targets’ opinions on whether the message was human or AI-written seemed influenced by their awareness of AI’s capability to use emojis and their previous encounters with AI using emojis. In our spear phishing messages, emojis appeared in 66% of messages written by AI and only 2% of messages written by humans.
5.4.3 Personalization
The message personalization was mentioned by 32% (8/25) of the targets. Five participants mentioned that a lack of personalization made them suspect the message was AI-generated. For instance, one target noted, “AI were the ones that were a little bit less personal" (T34), while another expressed doubt, saying, “There is no connection with me as a reader" (T41). Some targets believed the absence of personal pronouns or a personal introduction with the recipient’s name indicated AI authorship. One target observed, “All of these messages, but two, include my name" (T37), leading them to suspect those two were AI-generated. Conversely, four participants cited personalized messages as evidence of human authorship, providing similar reasons as those who suspected AI. This group of participants felt messages tailored to the recipient sounded human, with one target commenting, “it just sounds very personal" (T21). Additionally, recipients noted that human-sounding messages often began with an introduction of the sender, such as “this is Sarah from the [my university] student instructor program" (T41). Overall, personalized messages were deemed more human-like, while those lacking personalization were associated with AI.
5.4.4 Word Choice
The message word choice was mentioned by 24% (6/25) of the targets. Four of the targets claimed that the word choice sounded like AI, particularly the use of buzzwords and marketing words. One target explained that messages sound AI-generated when they use buzzwords that are associated with being “attention grabbers” because “AI would utilize those a lot” (T16). Another target described a message as “too specific” when it included words that humans “wouldn’t necessarily use” when describing instructional design (T18). A few of the targets agreed that some of the messages had sentences that a human would have worded differently (T09). Conversely, two targets gave evidence of human-sounding messages and both mentioned slang words. One target described “There’s so many slang terms on it, that it seems really human to me” (T34), while another defended the messages they thought to be human-created and explained, “oftentimes, it was just because I felt like the language seemed a little more slang-like" (T06). Messages with specific, marketing-like words were perceived as AI-written while messages with casual, slang words that sounded more conversational were perceived as human-written.
5.4.5 Content/Structure
The structure of the message content was mentioned by 24% (6/25) of the targets. Participants pointed out various aspects of structure that hinted at AI involvement, including repetitive phrases and/or awkward sentence construction. For instance, one participant thought a message was AI generated because ““it looks like there’s a template. So there’s like a flow that you know, there’s a pattern you see it as maybe started by a human” (T41). Another noted instances where messages seemed generic or mass-produced, akin to “the email that is sent to everyone in the school" (T41). One also pointed out inconsistencies within messages, such as repetitive phrases or ““two different things, in the same message” (T04). One target also noted that AI-generated messages tended to be more wordy, containing “lots of filler phrases" compared to human-authored messages (T37).
5.4.6 Grammar/Spelling
The grammar and spelling of the messages was mentioned by 24% (6/25) of the targets. Four of the targets explained that the grammar and spelling was ““too perfect” (T34), which led them to believe the message was AI generated. One described that “the grammar was almost too correct” (T06), while another described the messages as “pretty” (T34). Messages that were written “not how you would actually speak” were considered to be AI-generated (T13). Three of the targets gave evidence of spelling and grammar that seemed human-like. One target said “this one’s a legitimate person right there, because they spelled the ‘util source’ wrong” (T37), while another gave a similar reasoning and stated, “I don’t think that AI would have made the grammar mistake” (T21). If the messages included a “mix of good grammar with bad or some texting language” (T34), a human author became more believable. While the participants who noted grammar and spelling were split between AI and human-based evidence, most came to the same conclusion that AI would have perfect grammar and spelling while humans would have made errors.
5.4.7 Length of Message
The length of message was mentioned by 16% (4/25) of the targets. Participants assumed messages made by AI would be longer than human made messages. One mentioned that they would have “AI do the longer ones and I would write the shorter ones myself,” (T21) assuming those who created these messages would do the same. These four targets were correct in their hunch that AI messages would be longer. Across the messages, AI-generated messages have on average 41% higher character count per message, with 337.8 characters for AI, and 237.9 characters for human-authored messages.
5.4.8 URL
The message URL was mentioned by 8% (2/25) of the targets. Both of these targets explained that the included URL led them to believe the message was AI-generated. One target mentioned that the word “dot” was spelled out rather than typed so it “seemed like something maybe an AI would do” (T13). The other target concluded that altered links, such as bitly, are evidence of AI and stated that AI “changed the URL to be something a little bit easier to read and understand” because they have used ChatGPT before and remembered an output with a shortened link. Both targets concluded that a suspicious URL component leads to the conclusion that an AI message was generated. Most (71%) human authors did not include an actual URL, instead they included placeholders such as “[URL],” “site,” or “url.” In contrast, only one AI message had a placeholder. As discussed in the methods, all of these placeholders were replaced with a shortened link from tinyurl.
6 Discussion
Our discussion section is organized around our key research questions, followed by a reflection on the pros and cons of the TRAPD methodology.
6.1 AI vs. Human in Creating Persuasive Spear Phishing Messages
A growing body of literature has found evidence that AI-generated content from recent LLMs can outperform human-created content in various domains .[8, 29, 30, 31] For example, Zhang and Gosline compared content generated by GPT-4 with similar content created by professional content creators in the advertising field, finding that “Content generated by generative AI and augmented AI is perceived as of higher quality than that produced by human experts and augmented human experts" .[29] Nisbett and Spaiser explored the use of GPT-3 in creating moral statements supporting climate action, concluding that “GPT-3-generated statements are generally more convincing than human-generated statements" .[31] Similarly, in the domain of political speech, Palmer and Spirling found that “LLMs are capable of producing human-style arguments for different positions on subjects as varied as abortion, guns, immigration, and organ donation" and they could “out-perform human authors, though it varies by topic" .[30]
Fewer studies have examined the ability of LLMs to create malicious content tailored to individuals, such as we examine in this paper. Heiding et al. examined phishing messages created by humans (using the V-triad approach), GPT-4, both human+GPT-4 combined, and a control message from an existing phishing dataset .[32] They found that the human V-triad and the combined human V-triad and GPT-4 approaches led to the highest click-through rate in a field experiment, followed by GPT-4 and then the control group.[32] Although they did not personalize these messages to each individual, they did personalize them to a specific university context. They also provide a cost analysis that demonstrates how inexpensive the creation of spear phishing messages can be. Hazell also points out that LLM-generated spear phishing messages can be “realistic" and “cost-effective," but does not provide systematic evidence comparing humans to AI, such as presented in this paper .[8]
Our study contributes to this body of literature by comparing the effectiveness of AI-created and human-created spear phishing messages tailored to an individual target. We found that AI generally outperformed humans in creating spear phishing messages; however, the difference was not statistically significant, in part due to the relatively low sample size. As presented in Section 5.1.1, there is an 80% probability that GPT-4 is at least as good at or better than humans at creating these highly personalized spear phishing messages. While further research is justified in this area, we believe our findings provide a clear indication of the capabilities of LLMs for generating plausible spear phishing messages. It is important to note that these results are based on a very simple prompt, which could likely be improved. Furthermore, LLMs continue to improve in their quality, suggesting that AI will likely perform better at this task in the future. However, despite the fact that our human spear phishing authors had been trained on techniques to create spear phishing, it is possible that they could have been trained better. For example, they were not trained using the V-triad approach used so successfully by Heiding et al. .[32]
Participants in this study represented a range of job roles, including entry-level employees, mid-level managers, and executives across industries such as technology, healthcare, finance, and education. Despite this diversity, no significant differences were observed in the ability to differentiate AI-generated from human-generated messages, suggesting that the challenge of identifying these messages is consistent across demographic groups. However, qualitative responses indicated some role-specific variations in message evaluation. For instance, individuals in managerial positions found workplace-related messages more persuasive, while technical staff were more likely to scrutinize URLs and sender details. This reinforces the need for occupation-specific training to address diverse phishing vulnerabilities.
6.2 Characteristics of Convincing Spear Phishing Messages
Our statistical analysis found that targets considered job-related spear phishing messages twice as persuasive than messages related to hobbies or social media posts (38% intended click rate compared to 19% hobbies and 17% social media). This was true for both human and AI-generated messages. Spear phishing in the workplace carries significant risk, targeting both employees [33] and executives .[34] This suggests the need to be particularly vigilant for work-related spear phishing attacks. [35]
The qualitative findings from this study point to several phishing message features that make them more convincing than others. [9, 12] Our findings aligned with prior research, revealing that messages matching the receiver’s expectations in terms of the sender, context, and relevance of a message were more convincing; those with poor grammar and spelling were less believable; and messages relating to the recipient’s life were more persuasive .[28, 36]
We found that most of the features were either wholly persuasive or wholly unconvincing to all the targets. For example, personal relevance was completely correlated to the messages being persuasive while inaccuracies were completely correlated to the message being unconvincing. We also found that in some cases, some features may dissuade one person from clicking on a phishing message, while the same feature could influence another person to click. These findings were recently reported .[32] This suggests that some features may work better for certain individuals than others - a level of personalization that we have not yet examined, but seems possible to implement in future AI-generated spear phishing attacks that have more information about targets’ preferences. This personalization would be consistent with prior research. For example, one study found that people with different personality traits may be more susceptible to different types of attacks .[37]
A key focus of this paper was on personalization of messages since these were spear phishing messages tailored to individuals. Personalizing messages can lead to more persuasive messages [38, 10, 6], but can also raise red flags when the personalization is even slightly off. In this study, over two-thirds of participants were likely to believe messages that related to topics of personal interest and relevance, showing the power of personalized phishing messages when done well .[39, 40] However, two-thirds of participants also shared examples of messages that were not believed because they included content that was not personally relevant or included red flags, such as referencing organizations they had never heard of. Getting personalization right showed up in other categories as well. For example, many targets thought the technology communication medium (texting) was not appropriate for the type of message they received. Others recognized context inaccuracies, such as a colleague’s name that is made up. And while some people were more convinced by messages that had familiar senders, over two-thirds were dissuaded by the inclusion of senders that were suspicious and not someone the person would know. All of these findings suggest the need for further research not only on what makes phishing messages persuasive, but how persuasion techniques relate to personalization in spear phishing .[39] We believe future studies can build upon the categories and insights related to spear phishing messages identified in this paper.
6.3 AI vs. Human Message Identification
This study illustrated the limitations of humans in trying to identify AI-generated spear phishing messages. Our quantitative results show that targets guessed accurately only 52% of the time, where 50% would be expected from randomly guessing. Qualitative results suggest that many individuals had no idea how to even approach this task, as they struggled to know what criteria to use to differentiate them. This finding aligns with previous studies, such as those comparing AI-generated and human-written poetry, which demonstrate that people often struggle to differentiate between AI and human authorship .[41] The inability to reliably differentiate AI- from human-generated messages is significant because it suggests that traditional heuristics for identifying phishing (e.g., grammar or spelling errors) may no longer apply. As AI-generated content becomes more refined, training individuals to detect subtle markers of malicious intent—rather than relying on assumed human traits—will be critical.
Despite their lack of success, targets explained criteria they used when guessing which messages were AI generated. Their lack of confidence when assessing the source of messages suggests that their assessment criteria were based more on superficial or heuristic cues rather than hard and fast rules. Our results were consistent with previous research about human heuristics for AI-generated language .[42] In this study, participants noted that the presence of first-person pronouns, use of contractions, and references to family topics were a sign of a human-authored message. However, these cues were not reliable and interestingly, the AI-generated texts that were optimized for perceived humanity were often rated as “more human” than actual human-written ones. Our results showed a similar pattern; many targets noted that AI has a writing style that is "overly formal", or "less personal" even though it is not a consistent or reliable marker for AI authorship.
Still, it is useful to better understand the perceived limitations that people put on AI. Participants described a variety of criteria including specific linguistic features (the presence of emojis, how the URLs look, word choice), as well as stylistic choices (length of message, content structure, grammar, and spelling). They commonly noted that AI-generated messages frequently exhibit peculiar word choices, overly formal or inconsistent styles, and unnatural content structures. Moreover, AI-generated texts were often perceived as either too perfect or containing subtle but noticeable grammatical and spelling errors. Also, messages that lacked specific personal details or felt generic were more likely to be identified as AI-generated, since people assumed that AI could not create highly personal content. These perceptions are particularly problematic, since this study demonstrates that, in fact, AI is likely better at creating highly personalized phishing messages than humans.
These insights highlight the dual importance of distinguishing AI-generated messages: (1) to identify their unique features for detection and (2) to improve awareness about their potential misuse. In the long term, recognizing the origin of messages could inform defenses (human or automated) against these emerging threats.
6.4 The TRAPD Methodology
There are several pros of the TRAPD methodology. First, it provides an ethical way for personalized deceptive messages to be evaluated, since it starts with targets giving consent to share their own personal information in ways that can be “weaponized," while minimizing the effects of the deceptive messages. Second, the method provides data in a format that can statistically determine differences in the effectiveness of different types of deceptive messages (e.g., AI vs Human, topic of message). Third, it allows targets to provide qualitative insights about their decision-making when ranking messages. It is difficult to collect such data from other methods (e.g., field experiments where targets receive a single message in an authentic environment). Furthermore, grounding the discussion in the actual messages themselves (e.g., “why did you place this message in the most likely to click on position?") supports accurate self-reported reasoning. Having them do this for multiple messages in one sitting also helps them recognize patterns. Finally, this method can be a positive learning experience for the targets as they evaluate their own thinking and what is most and least likely to deceive them.
Survey methods are often criticized for lower ecological validity compared to field experiments. However, they remain effective for understanding participants’ perceptions and controlled comparisons. Prior research suggests, particularly in phishing studies, they remain valid in capturing behavioral variables, like participants’ perceptions of phishing messages[43]. Research also indicates that contextual factors can significantly influence participants’ behavior toward phishing attacks[44]. We strengthen TRAPD’s ecological validity by incorporating personalized messages tailored to participants’ own contexts, thus aligning their evaluations more closely with real-world decision-making. Although field studies are generally more ecologically valid, they can raise ethical concerns as studies of this nature often involve deception and getting permission to create personalized phishing messages without tipping off users by gaining consent can be difficult. TRAPD addresses both these concerns by using an adequately valid method of collecting data in a controlled, ethical environment.
As with all methods, there are also drawbacks. These drawbacks primarily come from the fact that the targets provide self-reported data in a lab environment. For example, it is possible that participants say they would click on a link in a message, when in fact they would not, or vice versa. Targets may have a social desirability bias, which would lead them to be less likely to admit falling for a deceptive message in the lab environment, if they recognize that the messages are deceptive in nature. Although it is hard to know how strong this tendency may be, the majority of our participants (80%) said they would click on the link in at least one message. Furthermore, those who said they were confident in their ability to identify phishing messages (in the questionnaire a month prior to the sorting/interview activity) said they would fall for fewer of them. It is also possible that individuals would act differently in more authentic contexts than what they expect when reporting in a lab. For example, a target may be less discerning when trying to quickly reply to a text message while riding a crowded bus than if they were at home on the couch. In our particular study, we used printouts of the text messages on the outline of a mobile phone to try and trigger similar thought patterns as they might have if looking on a real phone, but the lab environment and fact that it is paper and not digital was not possible to change. Additionally, this method focuses on the “content" of the messages, but cannot give input on some contextual factors, such as the source of the message. In the end, we have stronger confidence in the TRAPD method to identify differences between deceptive messages (and types of messages) than we do in the actual percentage of messages targets self-report as being willing to click on. Our observations of sessions with targets confirmed that they were fully engaged and took the sorting and threshold identification tasks very seriously.
On a practical level, we learned several things about effectively implementing the TRAPD methodology that we hope will be useful to others in the future. We originally had planned on having targets sort 16 items, but after pilot testing, decided to lower the number to 12. Our observations suggest that it took participants about 10-15 minutes to sort 12. We would not recommend going above that number. As expected, not all targets who started the study came back for their interview. We ended up with a 61% return rate (25/41) after multiple requests for targets to return for the sorting interview. We propose that future studies plan for similar or possibly lower return rates. It is critical to make clear to targets that they are consenting to come back for an interview (if at all possible), not just to complete an initial survey. In this particular study, the final target count of 25 was low enough that we did not get statistically significant results in some areas that we likely would have with a larger sample. Because we already solicited the human generated messages that are targeted to the original targets, it is not practical to increase the sample without having to redo essentially the entire study. Thus the importance of getting a high return rate when using human-created messages. One other limitation of the current study, which can inform future studies, was the need to handle URLs consistently. We found that the GPT-4 generally created made-up links within messages, while many humans just included text such as “[URL]" or “link" instead of an actual link. We used a tinyurl.com link in such cases. Future studies may consider either using the same link within all messages or enforcing the need for humans to create full URLs.
The TRAPD methodology can be used for a variety of contexts where personalized deceptive messages are created and evaluated. Indeed, an early study evaluating the effectiveness of a spear phishing training approach versus a control group used a less formally documented version of the TRAPD method [45]. We anticipate future studies using TRAPD to examine personalized disinformation, interactive AI-generated vishing attacks, and other forms of spear phishing. While using human-generated messages as part of TRAPD is clearly possible, there are advantages of using only automated messages. For example, imagine running a study that compares different LLMs or different prompts. The study could recruit targets, generate personalized deceptive content, and have targets rank order and describe their reasoning all within the same online survey. This is not possible with personalized human-generated content, due to the needed delay to create the personalized content, but would be possible with AI-only generated content. An online tool that supports TRAPD would likely help scale up participant numbers.
6.5 Implications of AI-enabled Spear Phishing
Currently, generative AI use is no longer a novelty among enthusiasts and researchers. As of 2024, a global survey by McKinsey & Company reported that 72% of organizations have adopted AI in at least one business function .[46] Despite the myriad of positive, legitimate uses that generative AI offers, there is increasing concern about the ethical challenges of its potential misuse. Particularly, there is an inherent risk that malicious actors can leverage these powerful models to be repurposed for spear phishing. With how effective LLMs can provide output through prompts that do not require much coding skill, this can potentially allow users who are not particularly technologically savvy to prompt AI to create spear phishing content. Accountability is also a big concern when dealing with AI abuse—should the developers also be accountable if their AI was used for nefarious uses? The language of existing laws regarding cybercrime are often aimed towards traditional phishing methods and may not be enough to address the unique challenges of AI.[47] A deeper exploration on the capabilities of AI and the safeguards required to prevent misuse will be a significant motivation for policymakers to promote discourse and implement standards.
Despite recent advancements in the field, there is still a significant gap in literature regarding AI-enabled spear phishing. Cybersecurity is an arms race between attackers and defenders, thus research to better understand these risks is essential. While many studies offer updated training methods to inform employees and end-users about phishing attacks [45], this study aligns with these efforts while also addressing the emerging threat of AI-generated spear phishing. While several studies have examined tools, frameworks, and AI or machine learning capabilities to detect phishing messages [48, 49, 50], little is known how publicly available LLMs can succeed in this endeavor. By evaluating how effectively publicly available LLMs can generate convincing spear phishing attacks, we can emphasize the need for countermeasures for potential malicious AI use. We are hopeful that this study will bridge the gap in the existing literature.
Understanding whether a phishing message was created by AI or a human has broader implications for cybersecurity policy. AI-generated messages are cheaper and faster to produce, but human-generated messages can introduce unique context and adaptability that AI might lack. Our findings suggest that distinguishing AI- from human-generated phishing messages is not merely an academic exercise but a practical necessity for developing tools that can detect and counteract evolving threats. Future studies should focus on refining this distinction and exploring how attackers might blend AI and human strategies to create even more convincing messages.
As the results from this study suggest, derivative studies may explore methods to enhance cybersecurity training using personalized spear phishing messages generated by AI. Tailoring training exercises to reflect realistic and specific phishing scenarios can improve user awareness and resilience. Future research could also investigate the psychological and behavioral factors that contribute to social engineering susceptibility. Additionally, collaborations between policymakers, AI developers, and cybersecurity experts are crucial to establish safeguards that mitigate AI misuse while preserving innovation. Research efforts could also focus on refining AI detection systems to better identify and flag AI-generated phishing attempts. Lastly, studies examining the long-term impact of these implications could provide actionable insights to shape global cybersecurity standards.
7 Limitations
This section describes the various limitations that exist in this pilot study:
7.1 Phishing Message Design
The phishing messages were crafted by students, which may not match the sophistication of the messages created by professional social engineers. Nevertheless, the messages were validated and reviewed by cybersecurity professors to ensure an acceptable quality of the human messages. Furthermore, for each prompt, three student-generated messages were evaluated, and the two most convincing were selected. The phishing messages deliberately lacked contextual cues, such as the senders’ email address or URL, to isolate and evaluate the effectiveness of the content itself, without the influence of other factors. This study would not factor in responses by participants when encountering suspicious senders or links. While it may affect the number of participants who would or would not click on a link, it is unlikely to affect the relative rankings of the messages based on their content.
7.2 Study Design
Our methodology used a self-reported approach to measure susceptibility, which may not accurately reflect actual behavior, as participants might downplay or overemphasize their susceptibility. This is not an uncommon methodology, as there are studies that use self reporting to assess convincingness of a message .[51, 52, 53] This method was also chosen to ensure the study remained ethical and avoided exposing participants to real spear phishing attacks or privacy violations. This method also allowed for direct comparison between messages which was useful to answer our research questions. Also, the relatively small sample size (25) limits the generalizability of the findings and statistical power of the study. However, as this is a pilot study, our primary purpose is to serve as a stepping stone for future research.
8 Conclusion
Deceptive messages that are personalized to a particular individual, such as spear phishing messages, can be highly persuasive. We are entering a new age, when LLMs can be used to generate personalized deceptive messages. This study used a novel methodology, we call TRAPD, to compare the efficacy of deceptive spear phishing SMS messages created by humans and AI. Although our statistical results are not definitive (due to a sample size of 25), we find a high likelihood (80%) that AI outperforms humans. We also find that messages related to jobs outperform those related to hobbies and items purchased online. This study highlights the capability of current-gen AI to replicate or even surpass humans in crafting convincing spear phishing messages.
Participants who were targeted with spear phishing messages described reasons they classified personalized smishing messages as particularly deceptive or easy to identify as fake. These were classified into the following key categories, in order of importance: characteristics of the URL, proper use of the technology communications medium of texting, use of inaccurate information, degree of personal relevance, application of the scarcity principle, messaging style, use of plausible rewards, and characteristics of the stated sender. These findings reinforce the importance of understanding the attributes of personalized deceptive messages, in order to develop countermeasures.
We also report on the failure of targets to identify which messages are created by AI versus humans. Targets were not confident in their ability to identify AI-generated content. They were often inconsistent in their assessment of the features that they thought indicated a message was AI-generated. Features that influenced their AI vs human message decisions included the message having emojis, characteristics of the message URL, word choice, message style, length of messages, structure of messages, grammar/spelling, and if and how personalization was used. This inability to reliably distinguish AI- from human-generated messages presents a critical challenge for cybersecurity. As AI-generated content becomes increasingly indistinguishable from human-authored material, individuals may rely on outdated heuristics, such as grammar errors, that are no longer applicable.
In summary, current LLMs can create highly deceptive spear phishing messages personalized to a target without targets having any idea that they are created by AI. Future research should focus on refining detection technologies and improving end-user training to address the evolving threat of AI-enabled spear phishing. By doing so, we can better anticipate and mitigate these emerging risks.